OpenStudy (anonymous):

the length of the hypotenuse of a right triangle is the square root of the sum of the squares of the measures of the other two legs of thr triangle. Approx. the length of the hypotenuse of a right triangle if the legs have measures 12 and 15

OpenStudy (anonymous):

19

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

at an accident scene, a police officer may determine the rate of speed, r, in miles per hour of the car by using the following fromula r= radical symbol 20l, where l is the length of the skid marks. How fast was a car going if the skid marks at the scene are 180 feet long?

OpenStudy (anonymous):

\[r =\sqrt{20 L}=\sqrt{20 *180}=60 \text{mph}\]